OpenStudy (anonymous):

If 8000 dollars is invested in a bank account at an interest rate of 5% per year, compounded continuously? How many years will it take for your balance to reach 20000 dollars?

OpenStudy (anonymous):

y=Ae^(kt)

OpenStudy (anonymous):

k=0.05 , the "growth constant" , which is 5%

OpenStudy (anonymous):

The initial value is 8000 , so A=8000 \

OpenStudy (anonymous):

y=8000e^(0.05t)

OpenStudy (anonymous):

20 000 = 8000e^(0.05t) Solve for t.

OpenStudy (anonymous):

e^(0.05t) = (20/8) 0.05t = ln(20/8) t= (1/0.05) ln(20/8)

OpenStudy (anonymous):

\[t= \frac{1}{0.05} \ln(\frac{20}{8})\] and that is in years. I don't have a calculator but you can sub into a calculator to get the actual value.

OpenStudy (anonymous):

" compounded continuously" means it is an exponential growth,

OpenStudy (anonymous):

thank you very much